# An Optimism-based Approach to Online Evaluation of Generative Models

### v1

@@ Abstract
Existing frameworks for evaluating and comparing generative models typically target an offline setting, where the evaluator has access to full batches of data produced by the models. However, in many practical scenarios, the goal is to identify the best model using the fewest generated samples to minimize the costs of querying data from the models. Such an online comparison is challenging with current offline assessment methods. In this work, we propose an online evaluation framework to find the generative model that maximizes a standard assessment score among a group of available models. Our method uses an optimism-based multi-armed bandit framework to identify the model producing data with the highest evaluation score, quantifying the quality and diversity of generated data. Specifically, we study the online assessment of generative models based on the Fréchet Inception Distance (FID) and Inception Score (IS) metrics and propose the FID-UCB and IS-UCB algorithms leveraging the upper confidence bound approach in online learning. We prove sub-linear regret bounds for these algorithms and present numerical results on standard image datasets, demonstrating their effectiveness in identifying the score-maximizing generative model.
@@ 1 Introduction
Deep generative models have achieved astonishing results across a wide array of machine learning datasets. Quantitative comparisons between generative models, trained using different methods and architectures, are commonly performed by evaluating assessment metrics such as Fréchet Inception Distance (FID) [17] and Inception Score (IS) [37]. Due to the growing applications of generative models to various learning tasks, the machine learning community has continuously adapted evaluation methodologies to better suit the characteristics of the newly introduced applications.
@@
In this work, we focus on the online evaluation of generative modeling schemes, where we consider a group of generative models and attempt to identify the model with the best score by assessing the fewest number of produced data. By limiting the number of generated samples in the assessment process, online evaluation can save on the costs associated with producing large batches of samples. Additionally, online evaluation can significantly reduce the time and computational expenses required to identify a well-performing model in assessing a large group of generative models.
@@
To measure the performance of an online evaluation algorithm, we use the regret notion widely-used in the online learning literature. In the online learning task, the evaluation algorithm selects a generative model and observes a mini-batch of samples generated by that model in each round. Our goal is to minimize the algorithm’s regret, defined as the cumulative difference between the scores of the selected models and the best possible score from the model set. Therefore, the learner’s regret quantifies the cost of generating samples from sub-optimal models. Our goal is to develop online learning algorithms that result in lower regret values in the assessment of generated data.
@@
To address the described online evaluation of generative models, we utilize the multi-armed bandit framework. In a standard multi-armed bandit problem, the online learner seeks to identify the arm with the highest expected value of a random score. While this formulation has been widely considered in many learning settings, it cannot solve the online evaluation of generative models, as standard IS and FID metrics do not simplify to the expectation of a random variable and represent a non-linear function of the data distribution.
@@
By deriving concentration bounds for the FID and IS scores, we propose the optimism-based FID-UCB and IS-UCB algorithms to address the online assessment of generative models. These algorithms apply the upper confidence bound (UCB) approach using data-dependent bounds that we establish for the FID and IS scores. Furthermore, we analytically bound the regret of the IS-UCB and FID-UCB methods, demonstrating their sub-linear regret growth assuming a full-rank covariance matrix of the embedded real data.
@@
We discuss the results of several numerical applications of FID-UCB and IS-UCB to standard image datasets and generative modeling frameworks. We compare the performance of these algorithms with the greedy algorithm, which selects the model with the highest estimated score, and a naive-UCB baseline that applies the UCB algorithm with a data-independent upper confidence bound. Our numerical results show a significant improvement in our proposed algorithms compared to the greedy and naive-UCB baselines. Additionally, we analyze the performance of FID-UCB with different image data embeddings, including InceptionNet.V3 [40], CLIP [7], and DINOv2 [33]. Our empirical results indicate satisfactory performance under these standard embeddings. The following summarizes the main contributions of this work:
@@
  • Proposing an online evaluation framework for generative models that aims to minimize the regret of misidentifying the score-maximizing model from online generated data.
@@
  • Developing the FID-UCB and IS-UCB algorithms by applying the upper-confidence-bound framework to our data-dependent estimation of the scores.
  • Proving sub-linear regret bounds for the proposed FID-UCB and IS-UCB algorithms.
@@
  • Demonstrating satisfactory empirical performance of FID-UCB and IS-UCB in comparison to the greedy and naive-UCB baseline algorithms.
@@ 2 Related Work
Assessment of deep generative models. The evaluation of generative models has been extensively studied in the literature. Several evaluation metrics are proposed, including distance-based metrics such as Wasserstein critic [3], Fréchet Inception Distance (FID) [17] and Kernel Inception Distance (KID) [5], and diversity/quality-based metrics such as Precision/Recall [36, 27], density and coverage [32], VENDI [13], and RKE [20]. In addition, the related works develop metrics quantifying the generalizability of the generative models, including the authenticity score [2], the FLD score [21], the Rarity score [16], and the KEN score [45] measuring the novelty of the generated samples. In this paper, we primarily focus on FID and Inception score, which have been frequently used for evaluating generative models.
@@
Role of embeddings in the quantitative evaluation results. Due to the high-dimensionality of images, evaluation of the generated images mostly relies on the embeddings extracted by pretrained networks on the ImageNet dataset, e.g., InceptionNet.V3 [40]. However, [32] shows that such pretrained-embeddings could exhibit unexpected behaviors. Recently, several large pretrained models have been proposed, including DINOv2 [33] and CLIP [7]. [39] shows that DINOv2-ViT-L/14 enables more interpretable evaluation of generative models. In addition, [28] demonstrates that FID scores computed based on the embedding extracted by CLIP agree more with human-based assessments. In this paper, we provide the numerical results for the mentioned embeddings extracted by different pretrained models.
@@
Online Learning using diversity-related evaluation metrics. Online learning is a sequential decision-making framework where an agent aims to minimize a cumulative loss function revealed to her sequentially. One popular setting is the multi-armed bandit (MAB), whose study dates back to the work of [29, 41]. At each step, the agent chooses among several arms, each associated with a reward distribution, and aims to maximize a pre-specified performance metric. The primary concern of this body of literature considers maximizing the expected return [1, 4, 6]. Recent works study performance metrics related to the variance or entropy of the reward distribution, including the mean-variance criterion [38, 47] in risk-sensitive MAB, and informational MAB (IMAB) where the agent maximizes the entropy rewards [44]. However, to the best of our knowledge, the evaluation of generative models has not been exclusively studied in an online learning context.
@@
Online training of generative models. Several related works focus on training generative adversarial networks (GANs) [14] using online learning frameworks. [15] proposes to train semi-shallow GANs using the Follow-the-Regularized-Leader (FTRL) approach. [10] shows that optimistic mirror decent (OMD) can be applied to address the limit cycling problem in training Wasserstein GANs (WGANs). Also, the recent paper [34] studies the no-regret behaviors of large language model (LLM) agents. This reference proposes an unsupervised training loss, whose minimization could automatically result in known no-regret learning algorithms. On the other hand, our focus is on the evaluation of generative models which does not concern the models’ training.
@@ 3.1 Inception Score
where $X_{\textup{g}}\sim p_{\textup{g}}$ is a generated image, $p_{Y|X_{\textup{g}}}$ is the conditional class distribution assigned by the InceptionNet.V3 pretrained on ImageNet [40], and $p_{Y_{\textup{g}}}=\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[p_{Y|X_{\textup{g}}}]$ is the marginal class distribution. Further, we have that $\textup{IS}=\exp[I(X_{\textup{g}};Y_{\textup{g}})]=\exp[H(Y_{\textup{g}})-H(Y_{\textup{g}}|X_{\textup{g}})]$, where $I(\cdot;\cdot)$ is the mutual information, and $H(\cdot)$ is the Shannon entropy. A higher IS implies that the images generated by the model have higher diversity, since $p_{Y_{\textup{g}}}$ would be more uniformly distributed to increase $H(Y_{\textup{g}})$, and possess higher fidelity, because $p_{Y|X_{\textup{g}}}$ is closer to a one-hot vector to enforce a smaller $H(Y_{\textup{g}}|X_{\textup{g}})$.
3.2 Fréchet Inception Distance
Fréchet Inception Distance (FID) is another standard metric for evaluating generative models. Let $f(x)\in{\mathbb{R}}^{d}$ denote the $d$-dimensional feature of an image $x$, which is extracted by the InceptionNet.V3 embedding. FID is the Fréchet distance [12] between the feature distributions of the generated images $p_{f(X_{\textup{g}})}$ and the real images $p_{f(X_{\textup{r}})}$, defined as
$\textup{FID}=\|\mu_{\textup{g}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{\textup{g}}+\Sigma_{\textup{r}}-2(\Sigma_{\textup{g}}\Sigma_{\textup{r}})^{\frac{1}{2}}]$ (2)
assuming that $p_{f(X_{\textup{g}})}={\mathcal{N}}(\mu_{\textup{g}},\Sigma_{\textup{g}})$ and $p_{f(X_{\textup{r}})}={\mathcal{N}}(\mu_{\textup{r}},\Sigma_{\textup{r}})$ are Gaussian distributions.
@@ 4 Online Evaluation of Generative Models
In this section, we introduce the framework of online evaluation of generative models, which is given in Protocol 1. We denote by ${\mathcal{G}}:=[G]$ the set of generative models. For each generator $g\in[G]$, we denote by $p_{g}\in\Delta({\mathcal{X}})$ its generative distribution over the space ${\mathcal{X}}$, which can be texts or images. Given an evaluation metric, e.g., FID or IS, the corresponding score of the generator $g\in[G]$ is denoted by $\nu_{g}$. The evaluation proceeds in $T$ steps. At each step $t\in[T]$, the evaluating algorithm ${\mathcal{A}}$ picks a generator $g_{t}\in[G]$ and collects a batch of generated samples $\{x_{t}^{j}\sim p_{g}\}_{j=1}^{b}$, where $b\in{\mathbb{N}}_{+}$ is the (fixed) batch size. The evaluator aims to minimize the regret
@@
Regarding the challenges of online evaluation of generative models, note that the empirical estimation of the score could be biased and generator-dependent. In addition, the generative distribution typically lies in a high-dimensional space, which makes it difficult to estimate the score from limited data. Moreover, the evaluation metrics often incorporate higher-order information of the generative distribution (e.g., computing FID requires the covariance matrix). Hence, analyzing the concentration properties of the score estimation would be challenging.
Protocol 1 Online Evaluation of Generative Models
@@
5 Online Evaluation based on Fréchet Inception Distance
In this section, we consider the online evaluation of generative models by Fréchet Inception Distance (FID). Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from model $g$, the empirical FID is computed by
$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}],$ (4)

### v3

# A Multi-Armed Bandit Approach to Online Selection and Evaluation of Generative Models

@@ Abstract
Existing frameworks for evaluating and comparing generative models consider an offline setting, where the evaluator has access to large batches of data produced by the models. However, in practical scenarios, the goal is often to identify and select the best model using the fewest possible generated samples to minimize the costs of querying data from the sub-optimal models. In this work, we propose an online evaluation and selection framework to find the generative model that maximizes a standard assessment score among a group of available models. We view the task as a multi-armed bandit (MAB) and propose upper confidence bound (UCB) bandit algorithms to identify the model producing data with the best evaluation score that quantifies the quality and diversity of generated data. Specifically, we develop the MAB-based selection of generative models considering the Fréchet Distance (FD) and Inception Score (IS) metrics, resulting in the FD-UCB and IS-UCB algorithms. We prove regret bounds for these algorithms and present numerical results on standard image datasets. Our empirical results suggest the efficacy of MAB approaches for the sample-efficient evaluation and selection of deep generative models. The project code is available at https://github.com/yannxiaoyanhu/dgm-online-eval.
@@ 1 INTRODUCTION
Deep generative models have achieved astonishing results across a wide array of machine learning datasets. Quantitative comparisons between generative models, trained using different methods and architectures, are commonly performed by evaluating assessment metrics such as Fréchet Inception Distance (FID) (Heusel et al.,, 2017; Stein et al.,, 2023) and Inception Score (IS) (Salimans et al.,, 2016). Due to the growing applications of generative models to various learning tasks, the machine learning community has continuously adapted evaluation methodologies to better suit the characteristics of the newly introduced applications.
@@
In this work, we focus on the online evaluation and selection of generative modeling schemes, where we consider a group of generative models and attempt to identify the model with the best score by assessing the fewest number of produced data. By limiting the number of generated samples in the assessment process, online evaluation can save on the costs associated with producing large batches of samples. Additionally, online evaluation can significantly reduce the time and computational expenses required to identify a well-performing model in assessing a large group of generative models. Figure 1 shows an example of FD-based evaluation of three pretrained models where the embeddings are extracted by InceptionNet.V3. While the offline evaluation procedure requires a large batch of generated data from each model, our online evaluation approach queries generated data from models adaptively and limits sample generation from the sub-optimal models, thus collecting samples from the best model more frequently (Figure 5).
Figure 1: FID-based evaluation and selection among CIFAR10 generative models: The standard offline evaluation requires a large batch of data from every model. In contrast, our proposed online approach leverages the UCB multi-armed bandit strategy to identify the best model with fewer generations from the suboptimal models.
@@
To measure the performance of an online evaluation algorithm, we use the regret notion widely used in the online learning literature (Bubeck and Cesa-Bianchi,, 2012). In the online learning task, the evaluation algorithm selects a generative model and observes a mini-batch of samples generated by that model in each round. Our goal is to minimize the algorithm’s regret, defined as the cumulative difference between the scores of the selected models and the best possible score from the model set. Therefore, the learner’s regret quantifies the cost of generating samples from sub-optimal models. We aim to develop online learning algorithms that result in lower regret values in the assessment of generated data.
@@
To address the described online evaluation of generative models, we utilize the multi-armed bandit (MAB) framework. In a standard MAB setup, the online learner seeks to identify the arm with the highest expected value of a random score. While this formulation has been widely considered in many learning settings, it cannot solve the online evaluation of generative models, as standard IS and FD metrics do not simplify to the expectation of a random variable and represent a non-linear function of the data distribution.
@@
By deriving concentration bounds for the FD and IS scores, we propose the optimism-based FD-UCB and IS-UCB algorithms to address the online assessment of generative models. These algorithms apply the upper confidence bound (UCB) approach using data-dependent bounds that we establish for the FD and IS scores. Furthermore, we analytically bound the regret of the IS-UCB and FD-UCB methods, demonstrating their sub-linear regret growth assuming a full-rank covariance matrix of the embedded real data.
@@
We discuss the results of several numerical applications of FD-UCB and IS-UCB to standard image datasets and generative modeling frameworks. We compare the performance of these algorithms with the Greedy algorithm, which selects the model with the highest estimated score, and a Naive-UCB baseline that applies the UCB algorithm with a data-independent upper confidence bound. Our numerical results show a significant improvement in our proposed algorithms compared to the Greedy and Naive-UCB baselines. Additionally, FD-UCB can attain satisfactory performance under the standard image data embeddings, including InceptionNet.V3 (Szegedy et al.,, 2016), DINOv2 (Oquab et al.,, 2024), and CLIP (Radford et al.,, 2021). Our work demonstrates the effectiveness of MAB-based approaches to the evaluation and selection of generative models. The following summarizes the main contributions of this work:
@@
  • Proposing a MAB-based evaluation framework for generative models that aims to minimize the regret of misidentifying the score-maximizing model from online generated data.
@@
  • Developing the FD-UCB and IS-UCB algorithms by applying the upper-confidence-bound framework to our data-dependent estimation of the scores.
  • Proving sub-linear regret bounds for the proposed FD-UCB and IS-UCB algorithms.
@@
  • Demonstrating satisfactory empirical performance of FD-UCB and IS-UCB in comparison to the Greedy and Naive-UCB baseline algorithms.
@@ 2 RELATED WORK
Assessment of deep generative models. The evaluation of generative models has been extensively studied in the literature. Several evaluation metrics are proposed, including distance-based metrics such as Wasserstein critic (Arjovsky et al.,, 2017), Fréchet Inception Distance (FID) (Heusel et al.,, 2017) and Kernel Inception Distance (KID) (Bińkowski et al.,, 2018), and diversity/quality-based metrics such as Precision/Recall (Sajjadi et al.,, 2018; Kynkäänniemi et al.,, 2019), density and coverage (Naeem et al.,, 2020), Vendi (Friedman and Dieng,, 2023), RKE (Jalali et al.,, 2023), and scalable FKEA-Vendi Ospanov et al., (2024). In addition, the related works develop metrics quantifying the generalizability of the generative models, including the authenticity score (Alaa et al.,, 2022), the FLD score (Jiralerspong et al.,, 2023), the Rarity score (Han et al.,, 2023), and the KEN score (Zhang et al., 2024b, ; Zhang et al., 2024a, ) measuring the novelty of the generated samples. In this paper, we primarily focus on FID and Inception Score, which have been frequently used for evaluating generative models.
@@
Role of embeddings in the quantitative evaluation results. Due to the high-dimensionality of images, evaluation of the generated images mostly relies on the embeddings extracted by pretrained networks on the ImageNet dataset, e.g., InceptionNet.V3 (Szegedy et al.,, 2016). However, Naeem et al., (2020) shows that such pretrained embeddings could exhibit unexpected behaviors. Recently, several large pretrained models have been proposed, including DINOv2 (Oquab et al.,, 2024) and CLIP (Cherti et al.,, 2023). Stein et al., (2023) shows that DINOv2-ViT-L/14 enables more interpretable evaluation of generative models. In addition, Kynkäänniemi et al., (2023) demonstrates that FID scores computed based on the embedding extracted by CLIP agree more with human-based assessments. In this paper, we provide the numerical results for the mentioned embeddings extracted by different pretrained models.
@@
Online learning using diversity-related evaluation metrics. Online learning is a sequential decision-making framework where an agent aims to minimize a cumulative loss function revealed to her sequentially. One popular setting is the multi-armed bandit (MAB), whose study dates back to the work of Lai and Robbins, (1985); Thompson, (1933). At each step, the agent chooses among several arms, each associated with a reward distribution, and aims to maximize a pre-specified performance metric. The primary concern of this body of literature considers maximizing the expected return (Agrawal,, 1995; Auer,, 2003; Bubeck and Cesa-Bianchi,, 2012). Recent works study performance metrics related to the variance or entropy of the reward distribution, including the mean-variance criterion (Sani et al.,, 2012; Zhu and Tan,, 2020) in risk-sensitive MAB, and informational MAB (IMAB) where the agent maximizes the entropy rewards (Weinberger and Yemini,, 2023). However, to the best of our knowledge, the evaluation of generative models has not been exclusively studied in an online learning context. In a concurrent work, Rezaei et al., (2025) aim to maximize kernel-based evaluation scores by finding a mixture of generative models. In this work, we primarily focus on identifying the best (single) model.
@@
Online training of generative models. Several related works focus on training generative adversarial networks (GANs) (Goodfellow et al.,, 2014) using online learning frameworks. Grnarova et al., (2018) proposes to train semi-shallow GANs using the Follow-the-Regularized-Leader (FTRL) approach. Daskalakis et al., (2018) shows that optimistic mirror decent (OMD) can be applied to address the limit cycling problem in training Wasserstein GANs (WGANs). Also, the recent paper Park et al., (2024) studies the no-regret behaviors of large language model (LLM) agents. This reference proposes an unsupervised training loss, whose minimization could automatically result in known no-regret learning algorithms. On the other hand, our focus is on the evaluation of generative models which does not concern the models’ training.
@@ 3.1 Inception Score
where $X_{\textup{g}}\sim p_{\textup{g}}$ is a generated image, $p_{Y|X_{\textup{g}}}$ is the conditional class distribution assigned by the InceptionNet.V3 pretrained on ImageNet (Szegedy et al.,, 2016), and $p_{Y_{\textup{g}}}=\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[p_{Y|X_{\textup{g}}}]$ is the marginal class distribution. Further, we have that $\textup{IS}=\exp[I(X_{\textup{g}};Y_{\textup{g}})]=\exp[H(Y_{\textup{g}})-H(Y_{\textup{g}}|X_{\textup{g}})]$, where $I(\cdot;\cdot)$ is the mutual information, and $H(\cdot)$ is the Shannon entropy. A higher IS implies that the images generated by the model have higher diversity, since $p_{Y_{\textup{g}}}$ would be more uniformly distributed to increase $H(Y_{\textup{g}})$, and possess higher fidelity, because $p_{Y|X_{\textup{g}}}$ is closer to a one-hot vector to enforce a smaller $H(Y_{\textup{g}}|X_{\textup{g}})$.
3.2 Fréchet Distance
Fréchet Distance (FD) (Dowson and Landau,, 1982) is another standard metric for evaluating generative models. Let $f(x)\in{\mathbb{R}}^{d}$ denote the $d$-dimensional feature of an image $x$. The FD between the feature distributions of the generated images $p_{f(X_{\textup{g}})}$ and the real images $p_{f(X_{\textup{r}})}$, which is computed by
$\textup{FD}=\|\mu_{\textup{g}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{\textup{g}}+\Sigma_{\textup{r}}-2(\Sigma_{\textup{g}}\Sigma_{\textup{r}})^{\frac{1}{2}}].$ (2)
The well-known FID metric (Heusel et al.,, 2017) is computed as the FD where the image data feature is extracted by InceptionNet.V3 (Szegedy et al.,, 2016). Recently, Kynkäänniemi et al., (2023); Stein et al., (2023) propose computing the FD distance using the CLIP (Radford et al.,, 2021) and DINOv2 (Oquab et al.,, 2024) embeddings, respectively, to boost the score’s ranking consistency with human evaluations.
@@ 4 ONLINE EVALUATION OF GENERATIVE MODELS
In this section, we introduce the framework of online evaluation of generative models, which is given in Protocol 1. We denote by ${\mathcal{G}}:=[G]$ the set of generative models. For each generator $g\in[G]$, we denote by $p_{g}\in\Delta({\mathcal{X}})$ its generative distribution over the space ${\mathcal{X}}$, which can be texts or images. Given an evaluation metric, e.g., FD or IS, the corresponding score of the generator $g\in[G]$ is denoted by $\nu_{g}$. The evaluation proceeds in $T$ steps. At each step $t\in[T]$, the evaluating algorithm ${\mathcal{A}}$ picks a generator $g_{t}\in[G]$ and collects a batch of generated samples $\{x_{t}^{j}\sim p_{g}\}_{j=1}^{b}$, where $b\in{\mathbb{N}}_{+}$ is the (fixed) batch size. The evaluator aims to minimize the regret
@@
Regarding the challenges of online evaluation of generative models, note that the empirical estimation of the score could be biased and generator-dependent. In addition, the generative distribution typically lies in a high-dimensional space, which makes it difficult to estimate the score from limited data. Moreover, the evaluation metrics often incorporate higher-order information of the generative distribution (e.g., computing FD requires the covariance matrix). Hence, analyzing the concentration properties of the score estimation would be challenging.
Protocol 1 Online Evaluation and Selection of Generative Models
@@
Use of regret metric. The choice of regret metric 3 follows the standard online learning literature for bandit problems (Bubeck and Cesa-Bianchi,, 2012). Particularly, if the evaluator can attain a sub-linear regret, then the overall selection converges to the best model with high probability. Additionally, in the online evaluation task, it is possible that the best model has a very similar performance to that of the second-best model. In such a case, the regret would take the difference between the models into account, and in a case where the top-$k$ models have similar scores, generating samples from them would weighted more equally. On the other hand, if the best two models have significantly different scores, the regret would sharply grow by picking the suboptimal models.
5 FRÉCHET DISTANCE-BASED ONLINE EVALUATION AND SELECTION
Avg. Regret
Figure 2: Online FD-based evaluation and selection among standard generative models: The $x$-axis is the number of online steps. At each step, the algorithm samples a batch of five generated images from the chosen model. The image data embeddings are extracted by CLIP (Cherti et al.,, 2023). Results are averaged over 20 trials.
[…]
In this section, we consider the online evaluation of generative models by Fréchet Distance (FD). Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from model $g$, the empirical FD is computed by
$\widetilde{\textup{FD}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}],$ (4)